# DemoTuner: Efficient DBMS Knobs Tuning via LLM-Assisted Demonstration Reinforcement Learning

### v1

@@ V-D1 Hint Priority-Aware PER
Prioritized experience replay [schaul2015per] is a key technique to achieve demonstration reinforcement learning. It assigns a priority to each experience (including both demonstrations and transitions) based on its significance, which affects its probability of being sampled for model updates. In detail, the probability of sampling a particular experience $i$ from the experience replay buffer is defined as:
@@ V-E Hint-Guided Reward Shaping
Reward is one of the most crucial components in RL. Reward shaping [ng1999rewardshaping] involves providing small intermediate synthetic rewards to the agent to expedite training convergence. In HA-DDPGfD, we design a hint-guided reward shaping technique to leverage the domain knowledge to accelerate the agent fine-tuning phase. Specifically, after obtaining the original reward $r_{t}$ defined as the performance improvement over default $perf_{0}$ (as described in Eq. (6a)), we adjust $r_{t}$ according to the divergence between the actual knob adjustments and the actions rec_actionrec\_action suggested by the tuning hints in current Match_HintsMatch\_Hints set. For instance, if a tuning hint suggests increasing the value of a specific knob but it actually decreases from action $a_{t-1}$ to action $a_{t}$, this discrepancy is treated as a violation of the tuning hint. It is worth noting that we only take the current matched tuning hints with priorities larger than a specified threshold into consideration. In detail, we use $\sum\mathbb{I}(a_{t},a_{t-1},h_{i})$ (where $h_{i}\in Match\_Hints$) to count the total number of violations, with $\mathbb{I}(\cdot)$ being the indicator function that takes the value of 1 when a violation occurs, and 0 otherwise. Then the shaping factor $f_{t}$ is obtained as:
@@ V-F The Proposed HA-DDPGfD Algorithm in DemoTuner
Then during the agent fine-tuning phase (Lines 14–27), DemoTuner iteratively interacts with the target DBMS via generating promising configurations and obtaining performance feedback to improve the policy. For each step, DemoTuner generates an action according to current policy network and then adjust it based on the hints’ priorities. After that, the adjusted action is evaluated and both 1-step and n-step transitions are stored in the experience replay buffer. For each episode, DemoTuner follows the proposed hpPER mechanism to select experiences to update the agent’s critic and actor networks. Specially, to better balance the short-term and long-term reward feedback, we also leverage a mix of 1-step and n-step returns to update the critic network as previous work [vecerik2017leveraging] did. Specifically, the n-step return is defined as follows:
@@ VI-A Experimental Platform
To evaluate the effectiveness and efficiency of DemoTuner, we conducted a series of experiments using two widely utilized database management systems MySQL (version 8.0.36) and PostgreSQL (version 16.1). For configuration evaluation, we deploy the target DBMS on a Linux server with an Intel processor of 8 physical cores, 32GB of RAM, and a 300GB SSD. The operating system is Ubuntu 22.04.2 with kernel version 6.5.0. In order to mitigate the potential performance interference, we make sure that nothing else is running on the server except the essential kernel processes and the target DBMS service when conducting performance tests to evaluate generated configurations. It is worth noting that DemoTuner currently supports two widely deployed DBMSs: MySQL and PostgreSQL. However, we can easily add supports for other DBMSs according to the following two steps. First, for tuning hints extraction, we need to collect source texts about knobs tuning for the new DBMS and change the target DBMS information in the prompt. Second, for offline agent training, we only need to modify the control script for automatic configuration evaluation of the new DBMS. The source code of DemoTuner is publicly available at [DemoTuner2025Codes] for further extension and improvement.
@@ VI-B Benchmark Applications
The main benchmark we employed for configuration evaluation is YCSB (Yahoo! Cloud Serving Benchmark) [cooper2010benchmarking], which is a frequently-used benchmarking tool for performance tests of DBMSs. Similar to WATuner [ge2021watuning], we also synthesize different workloads based on the proportions of operations (read, update, etc.) in YCSB. Specifically, we evaluate DemoTuner as well as the baseline methods with three different workload types including Read-Heavy (RH), Read-Write (RW), and Write-Heavy (WH), and list the concrete operation proportions in Table II. For each workload type, we generate 1,000,000 records for PostgreSQL and 500,000 records for MySQL, and each configuration evaluation involves 500,000 operations launched with 10 concurrency threads. The workload datasize and operation number are set according to the computing capability of our testbed as well as considering the time consumption of different experiments.
@@
In addition to YCSB, TPC-H (OLAP) and TPC-C (OLTP) workloads are also used to evaluate the adaptability of DemoTuner to unknown workloads in Section VII-B and provide a query-level performance improvement analysis in Section VII-E. According to GPTuner [lao2025gptuner], we set the scaling factor of TPC-H and TPC-C to be 1 and 200, respectively. Besides, the TPC-C benchmark operates with 32 terminals and a transaction rate of 50,000. These workload settings are implemented with BenchBase [difallah2013oltpbench].
@@ VI-C Selected Knobs and Tuning Hints
In our experiments, we use a set of search keywords such as PostgreSQL performance tuning and MySQL configuration tuning to collect source texts from various documents such as web forums, blog posts, official manuals and configuration files. GPT-3.5-Turbo [gpt2025] is utilized considering the tradeoff between costs and practical effectiveness. After tuning hints extraction, we prioritize the knobs according to their mentioned frequency since this metric well indicates the degree of correlation with performance. We also consult official configuration guides as well as empirical data to further adjust our selection. Following the above process, we finally select 19 performance-critical knobs for MySQL and 18 for PostgreSQL. Besides, we obtain 70 tuning hints (average 3.7 hints per knob) for MySQL and 104 (average 5.8 hints per knob) for PostgreSQL.
@@ VI-D Baseline Methods
To evaluate the performance of DemoTuner in DBMS knobs tuning, we select two recent related studies as the baselines which also tried to combine domain knowledge with machine learning methods, namely DB-BERT [trummer2023db] and GPTuner [lao2025gptuner]. Additionally, we also compare DemoTuner with CDBTune [zhang2019CDBTune], a representative method that uses deep reinforcement learning for knobs configuration. In our experiments, DemoTuner undergoes a total 200 iterations of configuration evaluation for agent pre-training and fine-tuning, with another 10 iterations conducted for online tuning. As follows, we give a brief introduction of each baseline and how we implement them.
  • DB-BERT [trummer2023db] is a recent NLP-enhanced online knobs tuning method which utilizes BERT to translate natural language hints into recommended settings. At runtime, it employs an iterative RL algorithm to aggregate, adapt, and prioritize these hints for guiding the selection of knobs settings to evaluate. We utilize the available implementation version from the authors and set the RL iteration number to be 210, which is equal to the total number of offline and online iterations of DemoTuner.
  • GPTuner [lao2025gptuner] first utilizes a LLM-based pipeline to collect and refine domain knowledge to achieve efficient knob selection and value range identification. It then explores the optimal knob settings using a coarse-to-fine Bayesian optimization framework, which sequentially explores two search spaces of different granularity. We use the authors’ version and leverage 200 initial sampling configurations (30 in the coarse manner and 170 in the fine manner) recommended by GPTuner to construct the initial Gaussian process model. The online tuning iteration number is also set to be 10.
  • CDBTune [zhang2019CDBTune] is an end-to-end automatic cloud database tuning system that employs deep deterministic policy gradient methods to find the optimal configuration in high-dimensional continuous spaces. We use the PyTorch library to build neural networks to implement DDPG in CDBTune. It also undergoes 200 iterations for agent offline training, with another 10 iterations conducted for online tuning.
@@ VII-E Query-level Performance Improvement Analysis
Like many previous workload-level DBMS knobs tuning studies, DemoTuner primarily targets improving the overall performance of a DBMS under typical workloads. In contrast, query-level knobs tuning works [li2019qtune, xiu2025hint, chen2025aqetuner] focus on optimizing parameters that affect the execution of individual queries (often high-cost or critical ones). To provide a fine-grained view of DemoTuner’s workload-level performance improvement, we analyze its per-query effects over the default configuration on PostgreSQL with TPC-H (22 different queries in total). As shown in Figure 9, DemoTuner improves the majority of queries (17/22), leading to the 23.23% workload-level performance improvement over default. These gains primarily stem from the changes in query plan and execution procedure induced by knobs tuning such as Reduced Disk I/O, Enhanced Memory Use, Altered Parallelism, etc. On the other hand, performance regressions of the other 5 queries reflect the inherent tension between workload-level and query-level configuration tuning: knobs tuned for the overall workload-level performance might push the query optimizer toward choices that degrade specific queries. Extending DemoTuner to perform workload-level knobs tuning while guaranteeing per-query performance is a promising direction for future work.
@@ VIII Conclusion
AI-Generated Content Acknowledgement
The authors confirm that no generative AI tools or technologies were used to create any content in this paper, including text, figures, and images. All content is entirely the result of the authors’ original work. While this article incorporates the use of a Large Language Model (GPT) as part of DemoTuner’s design and implementation process (detailed in Section IV-B), GPT was utilized solely as a research tool and was not employed in the generation of the paper’s content.

### v2

@@ V-D1 Hint Priority-Aware PER
Prioritized experience replay [28] is a key technique to achieve demonstration reinforcement learning. It assigns a priority to each experience (including both demonstrations and transitions) based on its significance, which affects its probability of being sampled for model updates. In detail, the probability of sampling a particular experience $i$ from the experience replay buffer is defined as:
@@ V-E Hint-Guided Reward Shaping
Reward is one of the most crucial components in RL. Reward shaping [26] involves providing small intermediate synthetic rewards to the agent to expedite training convergence. In HA-DDPGfD, we design a hint-guided reward shaping technique to leverage the domain knowledge to accelerate the agent fine-tuning phase. Specifically, after obtaining the original reward $r_{t}$ defined as the performance improvement over default $perf_{0}$ (as described in Eq. (6a)), we adjust $r_{t}$ according to the divergence between the actual knob adjustments and the actions rec_actionrec\_action suggested by the tuning hints in current Match_HintsMatch\_Hints set. For instance, if a tuning hint suggests increasing the value of a specific knob but it actually decreases from action $a_{t-1}$ to action $a_{t}$, this discrepancy is treated as a violation of the tuning hint. It is worth noting that we only take the current matched tuning hints with priorities larger than a specified threshold into consideration. In detail, we use $\sum\mathbb{I}(a_{t},a_{t-1},h_{i})$ (where $h_{i}\in Match\_Hints$) to count the total number of violations, with $\mathbb{I}(\cdot)$ being the indicator function that takes the value of 1 when a violation occurs, and 0 otherwise. Then the shaping factor $f_{t}$ is obtained as:
@@ V-F The Proposed HA-DDPGfD Algorithm in DemoTuner
Then during the agent fine-tuning phase (Lines 14–27), DemoTuner iteratively interacts with the target DBMS via generating promising configurations and obtaining performance feedback to improve the policy. For each step, DemoTuner generates an action according to current policy network and then adjust it based on the hints’ priorities. After that, the adjusted action is evaluated and both 1-step and n-step transitions are stored in the experience replay buffer. For each episode, DemoTuner follows the proposed hpPER mechanism to select experiences to update the agent’s critic and actor networks. Specially, to better balance the short-term and long-term reward feedback, we also leverage a mix of 1-step and n-step returns to update the critic network as previous work [35] did. Specifically, the n-step return is defined as follows:
@@ VI-A Experimental Platform
To evaluate the effectiveness and efficiency of DemoTuner, we conducted a series of experiments using two widely utilized database management systems MySQL (version 8.0.36) and PostgreSQL (version 16.1). For configuration evaluation, we deploy the target DBMS on a Linux server with an Intel processor of 8 physical cores, 32GB of RAM, and a 300GB SSD. The operating system is Ubuntu 22.04.2 with kernel version 6.5.0. In order to mitigate the potential performance interference, we make sure that nothing else is running on the server except the essential kernel processes and the target DBMS service when conducting performance tests to evaluate generated configurations. It is worth noting that DemoTuner currently supports two widely deployed DBMSs: MySQL and PostgreSQL. However, we can easily add supports for other DBMSs according to the following two steps. First, for tuning hints extraction, we need to collect source texts about knobs tuning for the new DBMS and change the target DBMS information in the prompt. Second, for offline agent training, we only need to modify the control script for automatic configuration evaluation of the new DBMS. The source code of DemoTuner is publicly available at [7] for further extension and improvement.
@@ VI-B Benchmark Applications
The main benchmark we employed for configuration evaluation is YCSB (Yahoo! Cloud Serving Benchmark) [5], which is a frequently-used benchmarking tool for performance tests of DBMSs. Similar to WATuner [12], we also synthesize different workloads based on the proportions of operations (read, update, etc.) in YCSB. Specifically, we evaluate DemoTuner as well as the baseline methods with three different workload types including Read-Heavy (RH), Read-Write (RW), and Write-Heavy (WH), and list the concrete operation proportions in Table II. For each workload type, we generate 1,000,000 records for PostgreSQL and 500,000 records for MySQL, and each configuration evaluation involves 500,000 operations launched with 10 concurrency threads. The workload datasize and operation number are set according to the computing capability of our testbed as well as considering the time consumption of different experiments.
@@
In addition to YCSB, TPC-H (OLAP) and TPC-C (OLTP) workloads are also used to evaluate the adaptability of DemoTuner to unknown workloads in Section VII-B and provide a query-level performance improvement analysis in Section VII-E. According to GPTuner [18], we set the scaling factor of TPC-H and TPC-C to be 1 and 200, respectively. Besides, the TPC-C benchmark operates with 32 terminals and a transaction rate of 50,000. These workload settings are implemented with BenchBase [8].
@@ VI-C Selected Knobs and Tuning Hints
In our experiments, we use a set of search keywords such as PostgreSQL performance tuning and MySQL configuration tuning to collect source texts from various documents such as web forums, blog posts, official manuals and configuration files. GPT-3.5-Turbo [27] is utilized considering the tradeoff between costs and practical effectiveness. After tuning hints extraction, we prioritize the knobs according to their mentioned frequency since this metric well indicates the degree of correlation with performance. We also consult official configuration guides as well as empirical data to further adjust our selection. Following the above process, we finally select 19 performance-critical knobs for MySQL and 18 for PostgreSQL. Besides, we obtain 70 tuning hints (average 3.7 hints per knob) for MySQL and 104 (average 5.8 hints per knob) for PostgreSQL.
@@ VI-D Baseline Methods
To evaluate the performance of DemoTuner in DBMS knobs tuning, we select two recent related studies as the baselines which also tried to combine domain knowledge with machine learning methods, namely DB-BERT [33] and GPTuner [18]. Additionally, we also compare DemoTuner with CDBTune [41], a representative method that uses deep reinforcement learning for knobs configuration. In our experiments, DemoTuner undergoes a total 200 iterations of configuration evaluation for agent pre-training and fine-tuning, with another 10 iterations conducted for online tuning. As follows, we give a brief introduction of each baseline and how we implement them.
  • DB-BERT [33] is a recent NLP-enhanced online knobs tuning method which utilizes BERT to translate natural language hints into recommended settings. At runtime, it employs an iterative RL algorithm to aggregate, adapt, and prioritize these hints for guiding the selection of knobs settings to evaluate. We utilize the available implementation version from the authors and set the RL iteration number to be 210, which is equal to the total number of offline and online iterations of DemoTuner.
  • GPTuner [18] first utilizes a LLM-based pipeline to collect and refine domain knowledge to achieve efficient knob selection and value range identification. It then explores the optimal knob settings using a coarse-to-fine Bayesian optimization framework, which sequentially explores two search spaces of different granularity. We use the authors’ version and leverage 200 initial sampling configurations (30 in the coarse manner and 170 in the fine manner) recommended by GPTuner to construct the initial Gaussian process model. The online tuning iteration number is also set to be 10.
  • CDBTune [41] is an end-to-end automatic cloud database tuning system that employs deep deterministic policy gradient methods to find the optimal configuration in high-dimensional continuous spaces. We use the PyTorch library to build neural networks to implement DDPG in CDBTune. It also undergoes 200 iterations for agent offline training, with another 10 iterations conducted for online tuning.
@@ VII-E Query-level Performance Improvement Analysis
Like many previous workload-level DBMS knobs tuning studies, DemoTuner primarily targets improving the overall performance of a DBMS under typical workloads. In contrast, query-level knobs tuning works [19, 39, 4] focus on optimizing parameters that affect the execution of individual queries (often high-cost or critical ones). To provide a fine-grained view of DemoTuner’s workload-level performance improvement, we analyze its per-query effects over the default configuration on PostgreSQL with TPC-H (22 different queries in total). As shown in Figure 9, DemoTuner improves the majority of queries (17/22), leading to the 23.23% workload-level performance improvement over default. These gains primarily stem from the changes in query plan and execution procedure induced by knobs tuning such as Reduced Disk I/O, Enhanced Memory Use, Altered Parallelism, etc. On the other hand, performance regressions of the other 5 queries reflect the inherent tension between workload-level and query-level configuration tuning: knobs tuned for the overall workload-level performance might push the query optimizer toward choices that degrade specific queries. Extending DemoTuner to perform workload-level knobs tuning while guaranteeing per-query performance is a promising direction for future work.